# Dynamical stability of evolutionarily stable strategy in asymmetric games

### v1

@@ II.2 Mutants in both or either subpopulations
While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS’s in Definitions 3b and 3c are, respectively, known as Nash–Pareto pair Hofbauer and Sigmund (1998) and two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature. Furthermore, it is obvious that when Definitions 1a, 1b, and 1c hold, the corresponding Definitions 3a, 3b and 3c hold as well; the converse need not hold good. However, Definitions 2a, 2b, and 2c neither implies and nor is implied by Definitions 3a, 3b and 3c, respectively.
@@ II.3 Equivalent definitions
As mentioned earlier each of aforementioned three sets has three equivalent definitions—versions (a), (b) and (c). Let us see the line of reasoning that establishes these equivalences. For illustrative purpose, let us focus on establishing the equivalence between Definitions 3a, 3b, and 3c; equivalence between Definitions 1a, 1b, and 1c, as well as the equivalence between Definitions 2a, 2b, and 2c, follows analogously.
First let us show that Definitions 3a and 3b imply each other. Let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be the ESS according to Definition 3a. Then for all infinitesimal $(\epsilon_{1},\epsilon_{2})\neq(0,0)$, Eq. (10) should be valid. Specifically, when $\epsilon_{1}=0$ and $\epsilon_{2}\neq 0$, then the first inequality in Eq. (10) does not contribute to the definition as an $\mathbf{x}$ not equal to $\mathbf{\hat{x}}$ does not exist; after all, $\epsilon_{1}=0$ means there is no mutant in subpopulation X. This implies that an equality, rather than a greater-than sign, gets imposed in the first condition of Eq. (10). Therefore, the second inequality must be satisfied, i.e., $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$. Moreover, this inequality is naturally also satisfied even when $\epsilon_{1}$ remains in the neighbourhood of zero. On the other hand for $\epsilon_{1}\neq 0$ and $\epsilon_{2}=0$, $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ when similar arguments are invoked. Hence, we can conclude the criteria of ESS for the resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ across the allowable ranges of $\epsilon_{1}$ and $\epsilon_{2}$ is the following:
@@
This conforms with Eq. (11). There is one case left to pondered upon: the mutant strategies $(\mathbf{{x}},\mathbf{{y}})\in\Sigma_{m}\times\Sigma_{n}$ such that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. Simultaneous existence of these equalities forces $\epsilon_{1}\neq 0$ and $\epsilon_{2}\neq 0$. Straightforward comparison of the coefficients of $\epsilon_{1}$ and $\epsilon_{2}$ in Eq. (10) tells us that the following conditions must be satisfied by $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ so as to be ESS as per Definition 3a:
$\text{If}~{}~{}\mathbf{y}\cdot{\sf C}\mathbf{x}>\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}~{}~{}\text{then}~{}~{}\mathbf{x}\cdot{\sf B}\mathbf{y}<\mathbf{\hat{x}}\cdot{{\sf B}}\mathbf{y}~{}~{}\text{and}~{}~{}\text{If}~{}~{}\mathbf{x}\cdot{\sf B}\mathbf{y}>\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}~{}~{}\text{then}~{}~{}\mathbf{y}\cdot{\sf C}\mathbf{x}<\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}.$ (15)
This shows that Definition 3a implies Definition 3b. Furthermore, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{1}$ and $\epsilon_{2}$, implies that whenever there is an ESS as per Definition 3b, it must be ESS as per Definition 3a as well. In conclusion, Definitions 3a and 3b are equivalent.
Next let us show that Definition 3a and 3c imply each other. Starting with Eq. (10), we multiply both sides of the first inequality of Eq. (10) by $\epsilon_{1}$ and add $(1-\epsilon_{1})\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})$ to both sides. Similarly, we multiply both sides of the second inequality of Eq. (10) by $\epsilon_{2}$ and add $(1-\epsilon_{2})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})$ to both sides. The two resultant inequality expressions can be rearranged to read:
$\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})>(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})$ (16a)
@@
We define $\mathbf{x_{1}}\equiv\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}$ and $\mathbf{y_{1}}\equiv\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}}$, and observe that ($\mathbf{x_{1}},\mathbf{y_{1}})$ lies within the infinitesimal neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ given that $\epsilon_{1}$ and $\epsilon_{2}$ are sufficiently small. Therefore, relabeling $(\mathbf{x}_{1},\mathbf{y}_{1})$ as $(\mathbf{x},\mathbf{y})$, we arrive at Eq. (13) exactly, thereby demonstrating that Definition 3a implies Definition 3c. Since these steps done to reach at this implication can be reversed, Definition 3c implies Definition 3a. Thus, we have established the equivalence between both these definitions. Moreover, in the light of the already established equivalence of Definitions 3a and Definition 3b, the equivalence of Definition 3b and Definition 3c is also automatically established.
III Dynamical stability of ESS in two-player–two-strategy bimatrix game
Among the three sets of definitions which one is the most useful one? While there might be various arguments in favour of each one, one desirable aspect of ESS is probably undeniable: ESS is a static game-theoretic concept—it basically is about existence of an equilibrium solution; how one reaches it starting from a scenario that is not in equilibrium requires invoking some sort of dynamics. In evolutionary game theory, replicator equation Taylor and Jonker (1978); Cressman and Tao (2014) is the most accepted paradigm in this respect. Thus, one expects the replicator equation’s fixed points that correspond to ESS to be asymptotically stable. For example, there is a claim Cressman and Tao (2014); Cressman (1996)—here adapted for bimatrix games—
Theorem 1: A 2ESS is a locally asymptotically stable fixed point of the following set of equations called replicator equations,
@@ III Dynamical stability of ESS in two-player–two-strategy bimatrix game
This theorem would make Definitions 3c (equivalently, 3a and 3b) most useful. However, let us consider an insightful example which we now present.
Let there be two populations and in each population there be two phenotypes. In population X there are two types L and M; in population Y there are two types G and H. Thus, the bimatrix game’s payoff elements is given as,
{blockarray}cccc&GH
{block}
c[ccc]L13
,2313,1M0,023,13
[…]
{block}
[…]
23131M002313
[…]
Evidently, there are two pure 2ESS: (L, G) and (M, H). This is seen by explicit checking the definition of 2ESS for them; below we show it for (L, G) [the check for (M, H) involves analogous steps]. If we denote the arbitrary mixed state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}x,1-x\end{pmatrix},\begin{pmatrix}y,1-y\end{pmatrix}\right)$, then the probability vector corresponding to the state (L, G) is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$. Because there are only two types, a short hand notation for arbitrary state is $(x,y)$ and for $(\hat{x},\hat{y})$ as $(1,1)$; furthermore, if $(x,y)$ is a neighborhood state of $(\hat{x},\hat{y})$, then one can take $x=1-\epsilon$ and $y=1-\delta$ where $\epsilon,\delta\rightarrow 0^{+}$. Therefore, we check the condition in Definition 3c (for 2ESS), to find $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=\epsilon(\frac{1}{3}-\frac{2}{3}\delta)$ which always greater than zero when $\epsilon$ and $\delta$ are sufficiently small. In other words, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}>\mathbf{x}\cdot{\sf B}\mathbf{{y}}$, implying that $(\hat{{x}},\hat{{y}})=(1,1)$ is 2ESS. We emphasize that due to presence of “or” in the definition, here checking only one of the inequalities is enough to establish 2ESS property.
Figure 1: Numerical demonstration illustrating various stability behaviours of 2ESS in different asymmetric interaction scenarios. All coloured markers represent 2ESS: green circles indicate locally asymptotically stable fixed points, red triangles indicate unstable fixed points, and the pink star represents a center. Subplot $(a)$ depicts the phase space trajectory of the replicator dynamics described by Eq. (19). It illustrates that the fixed point $(1,1)$, which is a 2ESS, is not stable—making it a counter example of Theorem 1—whereas the fixed point $(0,0)$, which meets the stricter conditions of Definitions 1c and 2c, is stable. Subplots $(b)$ and $(c)$ display the phase space trajectories of the replicator dynamics described by Eq. (26), using the payoff matrices from Eq. (32) and Eq. (33), respectively. Subplot $(b)$ provides a counterexample of Theorem $1^{\prime}$ for pure ESS, as the point $(1,1)$, despite being a 2ESS, is unstable. Subplot $(c)$ depicts an example where being a 2ESS can imply asymptotic stability of pure states, as $(0,0)$ is a 2ESS and asymptotically stable. Subplot $(d)$ represents the phase space trajectories of Eq. (26), with ${\sf A}={\sf D}={\sf 0}$, and ${\sf B}$ and ${\sf C}$ are payoff matrices of battle of sexes game Hofbauer and Sigmund (1998). Here, the interior fixed point, although a 2ESS, is a center. In Subplot $(e)$, ${\sf A}$ and ${\sf D}$ are taken from Eq. (32) (${\sf B}$ and ${\sf C}$ are same as in Subplot (d)), showing that introducing intra-specific interaction renders the fixed point, which is a mixed 2ESS, stable. This shows Theorem $1^{\prime}$ can be valid for mixed ESS only.
Now what about its dynamical stability when we see it as a fixed point of the corresponding replicator equation:
$\dot{x}$ $=$ $x(1-x)\left[\frac{1}{3}-\frac{2}{3}(1-y)\right],$ (19a)
$\dot{y}$ $=$ $-\frac{1}{3}y(1-y).$ (19b)
This is obtained by substituting the payoff matrices from Eq. (18) in Eq. (17). Clearly, the R.H.S. of Eq. (19b) is always negative, therefore $y$ always decreases to zero for any initial conditions, and $x$ either increases or decreases depending on the initial conditions. In fact, $(1,1)$ is an unstable rest point of replicator equations and $(0,0)$ is a stable fixed point (see Fig. 1(a)). The fact that 2ESS $(1,1)$ is not locally asymptotically stable fixed point contradicts Theorem 1.
Maybe with ESS of Definition 2c (equivalently, 2a and 2b) in place of 2ESS in Theorem 1 saves the theorem by not admitting (L,G) as ESS. We find that, at $(\hat{x},\hat{y})=(1,1)$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=\epsilon/3-\delta/3-2\epsilon\delta/3$ which is not greater than zero in all the neighbourhood. Thus, $(1,1)$ does not satisfy Definition 2c of ESS and not a stable fixed point. Let’s check Definition 2c for (M,H)—fixed point corresponding to which is stable (see Fig. 1(a)). We find that at $(\hat{x},\hat{y})=(0,0)$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=\epsilon/3+\delta/3-2\epsilon\delta/3$. We see that $(0,0)$ does satisfy Definition 2c of ESS and it is stable.
Finally, we turn to Definition 1c (equivalently, 1a and 1b). Noting that this condition of ESS is exactly that of 2ESS but with “or” replaced by “and”, for (L,G) to be ESS we need to show $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}>0$ in the neighbourhood of (L,G)—something we have already show above; and in addition, we have to show that $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}>0$. The LHS of latter inequality is easily shown to be $-\delta/3$, and hence the inequality is not obeyed by (L,G). In conclusion, with ESS of Definition 1c (like Definition 2c) in place of 2ESS in Theorem 1 saves the theorem by not admitting (L,G) as ESS. As final check, since the fixed point corresponding to (M,H) is asymptotically stable (see Fig. 1(a)), we hope that just like Definition 2c, Definition 1c also admits this as an ESS. This is can be easily verified: With $x=0+\epsilon\rightarrow 0$ and $y=0+\delta\rightarrow 0$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=\epsilon({1}/{3}-{2}\delta/3)>0$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=\delta/3>0$.
Above example shows that in some cases, Theorem 1, may be incorrect (see also Appendix A) but it can be rescued by redefining ESS to not mean 2ESS but that one implied by Definition 1c or Definition 2c. While the illustration involved corner fixed points, the validity of Theorem 1 is questionable even for 2ESS that correspond to interior fixed points. Let us elaborate in this in more general manner and for this purpose we take the equivalent—but physically more clear—route of relating ESS with strongly stable strategy Hofbauer and Sigmund (2003, 1998).
III.1 Pure Strategy

### v2

@@ II.2 Mutants in both or either subpopulations
While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS’s in Definitions 3b and 3c are, respectively, known as Nash–Pareto pair Hofbauer and Sigmund (1998, 2003) and two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature.
We notice that (6) directly matches with equilibrium conditions in Definition 3b. So obviously, ESS under or-definition implies NE; and strict NE trivially implies ESS. However, because of the presence of conjunction ‘or’ in Definition 3a, there arises an interesting point: It deceptively appears that a state that is not NE, e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ can in fact qualify as ESS—as an inspection of Definition 3a would suggest—because of the presence of conjunction ‘or’. However, this apparent anomaly is prevented since the definition allows for one of $\epsilon_{x}$ or $\epsilon_{y}$ to be zero independently. Specifically, in the case pointed out, mathematically setting $\epsilon_{x}=0$ (and $\epsilon_{y}\neq 0$), physically means that no mutants in X-subpopulation are allowed, i.e., no ${\bf x}\neq{\bf{\hat{x}}}$ can be used to satisfy $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, and hence the first inequality in (11) for infinitesimally small $\epsilon_{y}$. Owing to the conjunction ‘or’, the second inequality in (11) must be checked which anyway can’t be satisfied as we have assumed that $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. In conclusion, allowing $\epsilon_{x}$ and $\epsilon_{y}$ to vanish (but not simultaneously) prevents states that are not NE’s from qualifying as ESS’s. Unfortunately, this condition brings forth another issue: There cannot be any mixed ESS.
Consider Definition 3a version of or-definition. If the mutant state satisfies $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$, then (11) reduces to $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$. Consequently, setting any of the $\epsilon_{x}$ or $\epsilon_{y}$ to zero is not permissible, as it renders both conditions meaningless. For instance, if we set $\epsilon_{y}=0$, then both the comparisons $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ become invalid because $(\epsilon_{x},0)$ physically means there is no mutant (no choice of ${\bf y})$ in the Y-population. Thus, if you would want to allow mixed ESS in bimatrix games under or-definition, then we must adopt the convention that while looking for mixed ESS mutants have to be present in both X- and Y-subpopulations—this perspective may appositely be called the simultaneous deviation perspective. This view is adopted while defining the Nash–Pareto pair in Hofbuaer and Sigmund Hofbauer and Sigmund (1998). The alternate perspective—which we term face-value perspective is to permit all deviations outlined explicitly by the or-definition [as specified in Definition 3a, $0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$$0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $(\epsilon_{x},\epsilon_{y})\neq(0,0)$], regardless of the equalities: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. Under this perspective, there cannot be any mixed ESS in bimatrix games, as the ESS conditions would only be satisfied if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a strict NE. This view was adopted by Cressman Cressman and Tao (2014); Cressman (1992) while defining 2ESS. It is obvious that also for weak pure NE [e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$] the issue detailed above applies: Under simultaneous deviation perspective, it will be 2ESS but not so under face-value perspective.
@@ II.3 Equivalent definitions
As mentioned earlier each of aforementioned three sets (viz., and-, sum- and or-definitions) has three equivalent definitions—versions (a), (b) and (c). Let us see the line of reasoning that establishes these equivalences. For illustrative purpose, let us focus on establishing the equivalence between Definitions 3a, 3b, and 3c; equivalence between Definitions 1a, 1b, and 1c, as well as the equivalence between Definitions 2a, 2b, and 2c, follows analogously. Now, in what follows, we prove the assertion by demonstrating that Definitions 3a and 3b imply each other and Definitions 3a and 3c imply each other.
Now, let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be an ESS according to Definition 3a, so that for all infinitesimal $(\epsilon_{x},\epsilon_{y})\neq(0,0)$, inequalities in (11) are valid. First, we can realize that if either of the conditions $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ or $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}<\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ holds, then there always is some permissible value of $(\epsilon_{x},\epsilon_{y})$ for which the inequalities in (11) do not hold. For instance, if $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$, then for $\epsilon_{x}\neq 0$ and $\epsilon_{y}=0$, the first condition in (11) can clearly not be satisfied and the second condition becomes invalid because there is no mutant in Y-population hence the comparison in the second condition in (11) won’t make any physical sense. Hence, we can conclude that the plausible criteria—for resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$—across the allowable ranges of $\epsilon_{x}$ and $\epsilon_{y}$ to satisfy the inequalities in (11) is:
@@
Thus, four explicit possibilities arise from (15) as detailed below. We now show that if Definition 3a is satisfied under simultaneous deviation perspective, then one of the four cases (all corresponding to condition in (12)) must be satisfied with the natural requirement of conditions in (13) in the relevant case.
  1. 1.
    One can trivially see that either condition— $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$—is true when (11) holds good, or
  2. 2.
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
  3. 3.
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
  4. 4.
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ is true but in addition, since $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$, straightforward comparison of the coefficients of $\epsilon_{x}$ and $\epsilon_{y}$ in (11) tells us that the conditions in (13) must follow.
Above argument establishes that Definition 3a implies Definition 3b. Moreover, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{x}$ and $\epsilon_{y}$, implies that whenever there is an ESS as per Definition 3b, it must be ESS as per Definition 3a as well. In conclusion, Definitions 3a and 3b are equivalent. We remark, in passing, that this equivalence is intact under face-value perspective as well but then the only ESS’s possible are the ones that are strict NE.
Next let us show that Definitions 3a and 3c imply each other. Starting with Eq. (11), we multiply both sides of the first inequality of Eq. (11) by $\epsilon_{x}$ and add $(1-\epsilon_{x})\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ to both sides. Similarly, we multiply both sides of the second inequality of Eq. (11) by $\epsilon_{y}$ and add $(1-\epsilon_{y})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})$ to both sides. The two resultant inequality expressions can be rearranged to read:
$\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ (16a)
@@
We define $\mathbf{x_{1}}\equiv\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}$ and $\mathbf{y_{1}}\equiv\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}}$, and observe that ($\mathbf{x_{1}},\mathbf{y_{1}})$ lies within the infinitesimal neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ given that $\epsilon_{x}$ and $\epsilon_{y}$ are sufficiently small. Therefore, relabeling $(\mathbf{x}_{1},\mathbf{y}_{1})$ as $(\mathbf{x},\mathbf{y})$, we arrive at Eq. (14) exactly, thereby demonstrating that Definition 3a implies Definition 3c. Since these steps done to reach at this implication can be reversed, Definition 3c implies Definition 3a. Thus, we have established the equivalence between both these definitions. Moreover, in the light of the already established equivalence of Definitions 3a and Definition 3b, the equivalence of Definition 3b and Definition 3c is also automatically established.
It, however, is interesting to note that and-, sum-, and or-definitions are not equivalent and need not imply each other except that the and-definition implies the sum-definition for any ESS. This aspect of the definitions will be taken up in the next two sections.
III Bimatrix games
Among the three sets of definitions which one is the most useful one? Or each of definitions works equally well? While there might be various arguments in favour of each one, one desirable aspect of ESS is probably undeniable: ESS is a static game-theoretic concept—it basically is about existence of an equilibrium solution; how one reaches it starting from a scenario that is not in equilibrium requires invoking some sort of dynamics. In evolutionary game theory, replicator equation Taylor and Jonker (1978); Cressman and Tao (2014) is the most accepted paradigm in this respect. Thus, one expects the replicator equation’s fixed points that correspond to ESS to be asymptotically stable, that is, we expect the following claim to be satisfied an ESS Cressman and Tao (2014); Cressman (1996):
Theorem 1: An ESS is a locally asymptotically stable fixed point of the following set of equations called replicator equations,
@@ III Bimatrix games
It is well-known that NE corresponds to a fixed point of the replicator equation (17). Since we have already seen that under and-, sum-, and or-definitions ESS is NE, we effectively are sure that any ESS is a fixed point of the replicator equation.
We can categorize ESS into two types: pure ESS and mixed ESS. Pure ESS can represent either a strict or weak NE, while mixed ESS can only correspond to a weak NE. The nuances of stability arise specifically in cases where ESS is a weak pure NE. We will examine each category separately for bimatrix and asymmetric games. However, before proceeding, let us first identify which definition aligns with the defining concept of strong stability Hofbauer and Sigmund (1998, 2003).
III.1 Strong Stability